OpenStudy (anonymous):

determine how many times more intense the front rows of a rock concert(110 decibles) is compared to a large orchestra(98 decibles)

OpenStudy (lilg132):

110 = 98x 110 / 98 = 98x / 98 1.122448979591837 = x 1.122448979591837 is how many times the front rows of a rock concert is compared to a large orchestra

OpenStudy (lilg132):

dont forget to click good answer

OpenStudy (anonymous):

thats not right though

OpenStudy (lilg132):

sorry i think i may have done it the wrong way round?

OpenStudy (lilg132):

110x = 98 110x / 110 = 98 / 110 x = 0.8909090909090909 that is how many times the rock concert is compared to a large orchastera is that right?

OpenStudy (anonymous):

the answer says 15.85 times more intense

OpenStudy (lilg132):

were you given any other info for that question